# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 6.4 Two-Sided Matching and the Deferred Acceptance Mechanism
Their model includes two types of agents: students and colleges. What is unique about this model is that agents and resources are the same entities. Specifically, agents on one side of the market serve as resources for the other side, and vice versa.
Gale and Shapley considered two versions of their model. In the simpler one-to-one matching model, each agent has a unit demand. In the more general many-to-one matching model, agents on one side—students—each have unit demand, while agents on the other side—colleges—can have multi-unit demand. In this model, agents on one side have strict preferences over agents on the other side.
As an exercise in use-inspired basic theory, [^47] Gale and Shapley explicitly emphasize that some insights from their analysis can be useful in real life. Gale and Shapley focus on the existence of outcomes that satisfy the following criteria. [^48] Gale and Shapley (1962) considers a setting where each individual on one side is acceptable to every individual on the other. Here, we present the axiom in a more general context, without this assumption, as described in Roth and Sotomayor (1990).
@@
An outcome satisfies stability if no agent on either side is assigned an unacceptable partner, and no college-student pair strictly prefer each other over their current assignments, including nil assignments where a student is unmatched or a college has unfilled positions.
Through an iterative procedure called the individual-proposing Deferred Acceptance algorithm, widely regarded as one of the most important innovations in matching theory, Gale and Shapley proved that a stable outcome always exists. Moreover, they showed that each student weakly prefers their assignment under this procedure to any other stable outcome. Let’s refer to this outcome as the student-optimal stable matching.
The individual-proposing Deferred Acceptance algorithm proceeds in a series of steps as follows. [^49] This description is based on the more general cumulative offer process by Hatfield and Milgrom (2005). In the first phase of each step, each student “proposes” to their most-preferred acceptable college that has not yet rejected them. In the second phase, each college tentatively “holds” the most-preferred proposals it has received so far, up to its capacity, and rejects the rest. The algorithm terminates when no new proposals are made or rejected. At this point, the proposals on hold are finalized as assignments.
Gale and Shapley also introduced an analogous college-proposing Deferred Acceptance algorithm, where the roles in the algorithm are reversed. This algorithm generates an outcome that favors the colleges. We refer to this outcome as the college-optimal stable matching.
@@
Beyond this natural mapping between the two models, there is a deeper connection between the no justified envy axiom in student placement and the stability axiom in many-to-one matching—one that plays a central role in the “discovery-invention cycles” linking basic research in two-sided matching with practical advancements in school choice. To illustrate this connection, we first introduce the following mild efficiency axiom for student placement.
@@
Let the individual-proposing Deferred Acceptance mechanism, henceforth simply referred to as the Deferred Acceptance (DA) mechanism, be the direct student placement mechanism that selects the student-optimal stable matching of the induced many-to-one matching problem for any given student placement problem. [^50] Balinski and Sönmez (1999) and Abdulkadiroğlu and Sönmez (2003) refer to the same mechanism as the Gale-Shapley student-optimal mechanism and Gale-Shapley student-optimal stable mechanism respectively. Similarly, let the college-proposing Deferred Acceptance mechanism be the direct student placement mechanism that selects the college-optimal stable matching of the induced many-to-one matching problem for any given student placement problem.
@@
On the one hand, Theorem 6 provides yet another instance of external validation for Gale and Shapley’s celebrated invention of the institution-proposing Deferred Acceptance algorithm. On the other hand, together with Proposition 4, Theorems 5 and 6 immediately imply that, among all outcomes satisfying no justified envy, non-wastefulness, and individual rationality, MCSD produces the worst possible assignment for each student.
In retrospect, reliance on SSD is the root cause of MCSD’s shortcomings. To understand why, let us reassess its underlying iterative procedure. At each step, for each field-specific ranking, SSD allocates the seats of all colleges within that field. When SSD is applied, it operates as if the entire field is making coordinated offers to students, avoiding unnecessary proposals that students would decline due to more-preferred colleges within the same field. Thus, this procedure can be viewed as a “field-proposing” Deferred Acceptance algorithm that produces the student-pessimal stable outcome.
@@ 6.5 What Sets Deferred Acceptance Apart
Beyond helping clarify the root cause of the MCSD mechanism’s failure, the strong connection between the student placement and many-to-one matching models—especially Theorem 4—also positions the DA as a leading candidate for student placement. In fact, as established by the following result, the appeal of this mechanism is even greater than what is suggested by Theorem 4.
@@
In a system where no justified envy is essential, DA not only outperforms MCSD but also any other direct mechanism. Moreover, as can be seen from the following result, the edge of DA goes beyond welfare considerations.
@@
DA is the only mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and strategy-proofness.
Theorem 8 reveals that MCSD fails to satisfy strategy-proofness. This is also easy to see in a slight modification of Example 5, where either student could secure their first choice by reporting it as their only acceptable option, thereby improving upon their outcomes in the baseline scenario.
@@
Through a slight modification of Example 5, it becomes evident that MCSD does not respect priority improvements. Both students would receive their first choices if one of them were ranked last by all colleges in both fields, whereas in the baseline scenario, they receive their second choices. Thus, under MCSD, not only can a student benefit from misreporting their preferences, but they may also be penalized for improved performance on standardized tests.
In contrast, DA sets itself apart in this regard.
@@
DA is the only mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and respects priority improvements.
@@ 6.6 Failed Policy Attempt in Turkey and Lessons Learned
In 1997, even before submitting the paper for scholarly review, I shared my discovery with Turkish officials. Following several correspondences by mail and a meeting with the head of the centralized clearinghouse, ÖSYM,515151ÖSYM is the acronym for Öğrenci Seçme ve Yerleştirme Merkezi, which translates to English as the Student Selection and Placement Center. I received a formal letter from Ankara politely rejecting my proposal. [^52] See Figure 1 for a copy of the letter from the head of ÖSYM rejecting my proposal to replace MCSD with DA in Turkish college admissions, and Figure 2 for its English translation.
@@
My two main lessons from this experience were the following:
@@
    No matter how accurate my model and clean my analysis might be, my theoretical analysis alone will likely not cut it for my policy ambitions, and that I need to present more concrete value to the stakeholders in order to be able to influence policy.
@@
    How good the mechanism I advocate is unlikely to be very important for authorities with vested interests in maintaining the status quo , unless I also show that their current mechanism is really bad.
A few years later, these two lessons guided my interactions with the authorities at Boston Public Schools.
@@
Shortly after Balinski and Sönmez (1999) was published, I learned that, in the U.S., students were also being assigned seats at schools annually through various centralized clearinghouses—not for college admissions, but for K-12 admissions in several major school districts. This centralized assignment of (typically public) school seats to students was promoted as an alternative to neighborhood-based assignment and was commonly referred to as school choice.
@@
The second reason was more direct. Of all the student assignment mechanisms in the U.S. that we documented in Abdulkadiroğlu and Sönmez (2003) across more than a dozen school districts, not a single one satisfied no justified envy. Presumably, if no school district bothered to design a system that abides by this desideratum, it was likely considered potentially dispensable.
@@ 6.8 Top Trading Cycles: A Variant for School Choice
As presented earlier in Example 3, the incompatibility between no justified envy and Pareto efficiency arises in certain situations. There may be a student $i$ whose priority at a school $s$ is not high enough to secure a seat there, yet it is sufficient to “block” a Pareto-improving trade between other students. Suppose student $j$ has sufficiently high priority at school $s$, and student $k$ has sufficiently high priority at school $t$. If student $j$ prefers school $t$ over school $s$, and student $k$ prefers school $s$ over school $t$, then no justified envy prevents them from engaging in a Pareto-improving trade whenever student $i$ has higher priority at school $s$ than student $k$ [^54] This observation also forms the basis of the efficiency-adjusted deferred acceptance mechanism (EADAM) in Kesten (2010), along with related ideas in Ehlers and Morrill (2019) and Reny (2022). Thus, enforcing no justified envy comes at the cost of losing these types of Pareto-improving trades between priorities. While trading priorities earned via standardized tests may be implausible, why wouldn’t authorities allow such trades in applications where school priorities are determined by factors like proximity to home?
@@
This idea naturally connects school choice to the house allocation with existing tenants model (cf. Sections 5.45.6). In particular, two straightforward modifications—allowing for multiple identical positions at each school and school-specific priority rankings—naturally lead to an extension of the house allocation variant of the Top Trading Cycles (TTC) mechanism (see Section 5.6.1) to school choice.
@@
In the context of house allocation with existing tenants, TTC allows individuals to trade their priorities over collectively owned houses (vacant houses or those vacated during the algorithm) with their privately occupied houses. Since the priority order for each collectively owned house is identical in this model, there is no need for trading priorities between individuals.
@@
In contrast, in school choice, priorities differ across schools. Therefore, we adapted the TTC mechanism developed in Abdulkadiroğlu and Sönmez (1999) by removing the private-ownership objects and introducing heterogeneity in the priority orders of schools with multiple seats. The resulting school choice version of the Top Trading Cycles (TTC) mechanism trades students’ priorities whenever doing so leads to a Pareto improvement. In Abdulkadiroğlu and Sönmez (2003), we proposed this variant of TTC as an alternative to DA for environments where full enforcement of priorities can be relaxed.
@@
As a mechanism for school choice, TTC is similar to DA in that it satisfies individual rationality, non-wastefulness, strategy-proofness, and respect for priority improvements. However, TTC differs from DA by satisfying Pareto efficiency, but at the cost of losing no justified envy. According to Balinski and Sönmez (1999), DA is the clear choice in applications where no justified envy is indispensable. In Abdulkadiroğlu and Sönmez (2003), we proposed TTC as a main alternative for applications where no justified envy is desirable but a secondary consideration.
@@
At that time, we could not identify a single school district using a mechanism that satisfied no justified envy, which led us to consider TTC as a strong contender against DA for school choice in the U.S. However, time would show that we might have underestimated the appeal of the no justified envy axiom even in the U.S., and that DA would ultimately emerge as the clear winner between our two proposals.

### v4

@@ 6.4 Two-Sided Matching and the Deferred Acceptance Mechanism
Their model includes two types of agents: students and colleges. What is distinctive about this framework is that agents and resources are the same entities: agents on one side of the market serve as resources for those on the other, and vice versa.
Gale and Shapley present two versions of the model. In the simpler one-to-one matching model, each agent has unit demand. In the more general many-to-one matching model, agents on one side—students—each have unit demand, while agents on the other side—college—can have multi-unit demand. In this setting, agents on each side have strict preferences over agents on the other side.
As an exercise in a use-inspired basic theory, [^57] Gale and Shapley explicitly emphasize that some insights from their analysis can be useful in real life. Gale and Shapley focus on the existence of an outcome—called a matching—that satisfies the following axiom. [^58] Gale and Shapley (1962) consider a setting where every individual on one side is acceptable to every individual on the other, whereas we follow Roth and Sotomayor (1990) in presenting the axiom in a more general context without this assumption.
@@
A matching satisfies stability if no agent on either side is assigned an unacceptable partner, and no college-student pair strictly prefer each other over their current assignments, including nil assignments where a student is unmatched or a college has unfilled positions.
One of the most important innovations in matching theory is an iterative procedure that guarantees the existence of a stable matching.
Individual-Proposing Deferred Acceptance Algorithm.
The algorithm proceeds in a series of steps, each consisting of two phases. In the first phase of each step, each student “proposes” to their most-preferred acceptable college that has not yet rejected them. In the second phase, each college tentatively “holds” the most-preferred acceptable proposals it has received in that step, up to its capacity, and rejects the rest. The algorithm terminates when no new proposals are made or rejected. At this point, the proposals on hold are finalized as assignments.
We next illustrate this algorithm with an example, depicted in Figure 10.
Example 7.
There are four students, Alp, Banu, Cora, and Diya, and three schools: $X$, $Y$, and $Z$, where school $X$ has two seats and schools $Y$ and $Z$ each have one. The preferences of the students and schools over each other are as follows:
[…]
Step 1. Alp, Banu, and Cora propose to their top-choice school, $X$, while Diya proposes to her top choice, $Y$. School $X$ tentatively holds the two best proposals—from Alp and Cora—and rejects Banu. School $Y$ tentatively holds its only proposal from Diya.
Step 2. Having been rejected by $X$, Banu proposes to her second choice, $Y$. Alp, Cora, and Diya maintain their offers from Step 1. Because its capacity is not exceeded this step, school $X$ continues to hold the proposals from Alp and Cora. Of the two proposals received by $Y$, it tentatively holds the better one, from Banu, and rejects Diya.
Step 3. Having been rejected by $Y$, Diya proposes to her second choice, $X$. Alp, Banu, and Cora maintain their offers from Step 2. School $X$ now holds the two best proposals—from Diya and Alp—and rejects Cora. School $Y$ tentatively holds its only proposal this step, from Banu.
Step 4. Having been rejected by $X$, Cora proposes to her next choice, $Z$. Alp, Banu, and Diya maintain their offers from Step 3. Since no school receives more proposals than its capacity, no student is rejected in this step, and all offers are finalized.
The individual-proposing deferred acceptance algorithm concludes with Alp and Diya each receiving a seat at $X$, Banu receiving a seat at $Y$, and Cora receiving a seat at $Z$.
Figure 10: Execution of the individual-proposing deferred acceptance algorithm in Example 7. The top panel illustrates the setting with student preferences and school rankings. At each step, every student makes a proposal to their most preferred school that has not rejected them in previous steps, as indicated by a boxed proposal. Each school then reviews all proposals received in that step (also shown with boxes) and rejects the lowest-ranked proposals exceeding its capacity (2 for $X$, 1 each for $Y$ and $Z$). Rejections in the current step are marked by a line crossing the proposal box, while rejections from previous steps appear in faint gray with their crossed lines retained.
In their celebrated paper, Gale and Shapley proved that the individual-proposing deferred acceptance algorithm always yields a stable matching. They also showed that each student weakly prefers their assignment under this matching to their assignment under any stable matching, thereby called the individual-optimal stable matching (or student-optimal stable matching in our setting). They also introduced the analogous college-proposing deferred acceptance algorithm, in which the roles of students and colleges are reversed. This version yields a stable matching that favors the colleges, known as the college-optimal stable matching.
For illustration, the reader can easily verify that in Example 7, this college-proposing version of the algorithm terminates in two steps, assigning seats at $X$ to Alp and Diya—as in the individual-proposing version—but seats at their third-choice schools, $Z$ and $Y$, to Banu and Cora, in contrast to their second choices assigned under the individual-proposing version.
@@
Beyond this natural mapping between the two models, there is a deeper connection between the no justified envy axiom in student placement and the stability axiom in many-to-one matching—one that plays a central role in the “discovery–invention cycles” linking basic research in two-sided matching with practical advancements in school choice (see Figure 36). To illustrate this connection, we first introduce the following mild efficiency axiom for student placement.
@@
Let the individual-proposing deferred acceptance mechanism, henceforth simply referred to as the deferred acceptance mechanism (DA), be the direct student placement mechanism that selects the student-optimal stable matching of the induced many-to-one matching problem for any given student placement problem. [^59] Balinski and Sönmez (1999) and Abdulkadiroğlu and Sönmez (2003) refer to the same mechanism as the Gale-Shapley student-optimal mechanism and Gale-Shapley student-optimal stable mechanism respectively. Similarly, let the college-proposing deferred acceptance mechanism be the direct student placement mechanism that selects the college-optimal stable matching of the induced many-to-one matching problem for any given student placement problem.
@@
On the one hand, Theorem 6 provides yet another instance of external validity for Gale and Shapley’s celebrated invention of the college-proposing deferred acceptance algorithm. On the other hand, together with Proposition 4, Theorems 5 and 6 immediately imply that, among all outcomes satisfying no justified envy, non-wastefulness, and individual rationality, MCSD produces the worst possible assignment for each student.
In retrospect, reliance on SSD is the root cause of MCSD’s shortcomings. To see why, we reassess its underlying iterative procedure. At each step, within each field-specific ranking, SSD allocates the seats of all colleges in that field. When applied, SSD operates as if the entire field makes coordinated offers to students, avoiding unnecessary proposals that students would decline in favor of more-preferred colleges within the same field. In this sense, the procedure functions as a “field-proposing” deferred acceptance algorithm that yields the student-pessimal stable outcome.
@@ 6.5 What Sets Deferred Acceptance Apart
Beyond clarifying the root cause of the MCSD mechanism’s failure, the strong connection between the student placement and many-to-one matching models—especially Theorem 4—also positions DA as a leading candidate for student placement. Indeed, as the following result establishes, the appeal of this mechanism on welfare grounds is even stronger than Theorem 4 alone suggests in settings where no justified envy is indispensable.
@@
In other words, in a system where no justified envy is essential, DA not only outperforms MCSD but also any other direct mechanism. Moreover, as the next result shows, the advantage of DA extends beyond welfare considerations.
Figure 11: MCSD neither satisfies strategy-proofness nor does it respect priority improvements. The left and right panels represent two modifications of Example 6, depicted in Figure 9. The left panel illustrates a case where student A misrepresents their true preferences by removing their second-choice school $Y$ and reporting only their first-choice school $X$. The right panel shows a case where student A loses priority in Science, now ranked second in both categories. In both scenarios, student B is tentatively assigned seats at both schools in Step 1 and removes their second choice. In Step 2, both students receive tentative assignments at their first choices only, and the procedure terminates. Since student A receives their second choice under MCSD in Example 6, they benefit either by manipulating their preferences (left panel) or by losing priority (right panel).
@@
DA is the unique mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and strategy-proofness.
Observe that Theorem 8 implies that MCSD does not satisfy strategy-proofness. As illustrated in the left panel of Figure 11, this is also evident in a slight modification of Example 6, where either student can single-handedly secure their first choice by reporting it as their only acceptable option, thereby improving upon their outcome in the baseline scenario.
@@
Through another slight modification of Example 6, it also becomes evident that MCSD does not respect priority improvements. As illustrated in the right panel of Figure 11, both students would receive their first choices if one of them were ranked last by all colleges in both fields, whereas in the baseline scenario they receive their second choices. Thus, under MCSD, a student may not only benefit from misreporting preferences but also be penalized for improved performance on standardized tests.
In contrast, DA sets itself apart in this regard too.
@@
DA is the unique mechanism that satisfies individual rationality, non-wastefulness, no justified envy, and respects priority improvements.
@@ 6.6 Failed Policy Attempt in Turkey and Lessons Learned
In 1997, even before submitting the paper for scholarly review, I shared my discovery with Turkish officials. Following several correspondences by mail and a meeting with the head of the centralized clearinghouse, ÖSYM,606060ÖSYM is the acronym for Öğrenci Seçme ve Yerleştirme Merkezi, which translates to English as the Student Selection and Placement Center. I received a formal letter from Ankara politely rejecting my proposal. [^61] See Figure A1 for a copy of the letter from the head of ÖSYM rejecting my proposal to replace MCSD with DA in Turkish college admissions, and Figure A2 for its English translation, both in the Supplemental Materials.
@@
This was an eye-opening experience, leaving me with three inescapable takeaways:
@@
    What matters to authorities may differ from what matters to academic economists.
@@
    How good the mechanism I advocate is unlikely to matter much for authorities with vested interests in maintaining the status quo, unless I also show that their current mechanism is really bad.
  3. 3.
    No matter how accurate my model and clean my analysis might be, my theoretical analysis alone will likely not cut it for my policy aspirations. I need to present more concrete value to stakeholders in order to influence policy.
A few years later, these takeaways guided my interactions with the authorities at Boston Public Schools.
@@
Shortly after Balinski and Sönmez (1999) was published, I learned that, in the U.S., students were also being assigned seats at schools annually through various centralized clearinghouses—not for college admissions, but for K–12 admissions in several major school districts. This centralized assignment of (typically public) school seats to students was promoted as an alternative to neighborhood-based assignment and was commonly referred to as school choice.
@@
The second reason was more direct. Of all the student assignment mechanisms in the U.S. that we documented in Abdulkadiroğlu and Sönmez (2003) across more than a dozen school districts, not a single one satisfied no justified envy. Presumably, if no school district bothered to design a system that abides by this desideratum, it couldn’t have been indispensable.
@@ 6.8 Top Trading Cycles: A Variant for School Choice
As presented earlier in Example 4, there is, in general, an incompatibility between no justified envy and Pareto efficiency. There may be a student $i$ whose priority at a school $s$ is not high enough to secure a seat there, yet still sufficient to block a Pareto-improving trade between other students. Suppose student $j$ has sufficiently high priority at school $s$, and student $k$ has sufficiently high priority at school $t$. If student $j$ prefers school $t$ over school $s$, and student $k$ prefers school $s$ over school $t$, then no justified envy prevents them from engaging in such a Pareto-improving trade whenever student $i$ has higher priority at school $s$ than student kk. [^63] This observation also forms the basis of the efficiency-adjusted deferred acceptance mechanism (EADAM) in Kesten (2010), along with related ideas in Ehlers and Morrill (2019) and Reny (2022).
Thus, enforcing no justified envy comes at the cost of forgoing Pareto-improving trades of priorities. While trading priorities earned via standardized tests may be implausible, why would authorities not allow such trades in applications where school priorities are determined by factors like proximity to home or even a lottery draw?
@@
This idea naturally connects school choice to the house allocation with existing tenants model (cf. Sections 5.45.6). In particular, two straightforward modifications to the model—(1) allowing multiple identical positions at each school and (2) introducing school-specific priority rankings—lead naturally to an extension of the HA–TTC algorithm (see Section 5.6.1) for school choice.
@@
In the house allocation with existing tenants model, HA–TTC not only permits trades of privately owned (earmarked) houses by existing tenants, but also allows individuals to trade their priorities over central-planner-owned houses (vacant or vacated during the algorithm) for these privately owned houses. Since each central-planner-owned house has the same priority order, there is no need to trade priorities across individuals.
@@
In contrast, in school choice, student priorities differ across schools. We therefore adapted the HA–TTC algorithm by removing private-ownership objects and introducing heterogeneous priority orders across schools with multiple seats. The resulting algorithm, the School Choice variant of Top Trading Cycles (SC–TTC), trades priorities whenever this yields a Pareto improvement.
Using the same setting as in Example 7, we next illustrate the SC–TTC algorithm, depicted in Figure 12.
Example 8.
There are four students, Alp, Banu, Cora, and Diya, and three schools: $X$, $Y$, and $Z$, where school $X$ has two seats and schools $Y$ and $Z$ each have one. The preferences of the students and schools over each other are as follows:
[…]
At each step, each student points to their most-preferred school with available seats, and each school points to its highest-priority student still unassigned.
In Step 1, a cycle forms among Alp, $X$, Diya, and $Y$. Alp and Diya leave the process with seats at their first-choice schools, $X$ and $Y$, respectively. School $Y$ also leaves the process with all its seats filled, while $X$ remains with one unfilled seat.
In Step 2, a cycle forms between Cora and $X$. Cora leaves the process assigned to the remaining seat at $X$, and $X$ exits the process.
In Step 3, a cycle forms between Banu and $Z$. Banu leaves the process with a seat at $Z$, which then exits the process, terminating the algorithm.
SC–TTC concludes with Alp and Cora each receiving a seat at $X$, Diya receiving a seat at $Y$, and Banu receiving a seat at $Z$.
Figure 12: Execution of the SC–TTC algorithm in Example 8. The top-left panel indicates the setting. In Steps 1–3, the smaller nodes represent students, and the larger nodes represent schools and their unfilled seats. At each step, students point to their highest-ranked school with unfilled seats, and each school points to its highest-priority student who is still unassigned. When a cycle forms, indicated by boldface arrows, each participating student leaves the process with a seat at the school they point to. In Steps 2 and 3, students who have left the process with their assignments, and schools with no remaining seats, are shown as faint gray nodes.
It is instructive to observe that, under the outcome of SC–TTC in Example 8, not only does Banu prefer Diya’s assignment, $Y$, to her own, $Z$, which is her last-choice school, but she also has higher priority at school $Y$ than Diya. That is, Banu has justified envy toward Diya. This becomes possible under SC–TTC because Alp, who has higher priority than Banu at $Y$, trades his priority with Diya in exchange for her priority at school $X$, his first choice.
@@
As a mechanism for school choice, SC–TTC is similar to DA in that it satisfies individual rationality, non-wastefulness, strategy-proofness, and respect for priority improvements. However, SC–TTC differs from DA by satisfying Pareto efficiency, but at the cost of potentially violating no justified envy. As established by Balinski and Sönmez (1999), DA is the clear choice in applications where no justified envy is indispensable. By contrast, Abdulkadiroğlu and Sönmez (2003) proposed SC–TTC as the main alternative for applications where no justified envy, though desirable, is secondary and can therefore be relaxed.
@@
At that time, we could not identify a single school district using a mechanism that satisfied no justified envy, which led us to consider SC–TTC as a strong contender against DA for school choice in the U.S. However, time would show that we might have underestimated the appeal of the no justified envy axiom even in the U.S., and that DA would ultimately emerge as the clear winner between our two proposals.